The rate constant of a chemical reaction increased from 0.100 s−1 to 3.00 s−1 upon raising the temperature from 25.0 ∘C to 37.0 ∘C .

Please Calculate the value of (1/T2−1/T1) where T1 is the initial temperature and T2 is the final temperature.
1
Expert's answer
2020-06-05T07:28:47-0400

The rate "k" of a chemical reaction dependence on the absolute temperature "T" is described by Arrhenius law:

"k = A·e^{-\frac{E_a}{RT}}" ,

where "A" is the pre-exponential factor, "E_a" is the activation energy and "R" is the universal gas constant.

The two measurements of the rate constants at two different temperatures are enough to calculate the values of the pre-exponential factor and of the activation energy:

"k_1 = A·e^{-\frac{E_a}{RT_1}}"

"k_2 = A·e^{-\frac{E_a}{RT_2}}"

"\text{ln}(\frac{k_1}{k_2}) = \frac{E_a}{R}(\frac{1}{T_2} - \frac{1}{T_1})"

"E_a = R·\frac{\text{ln}(\frac{k_1}{k_2})}{\frac{1}{T_2} - \frac{1}{T_1}}" .

The value of (1/T2−1/T1) is:

"\frac{1}{T_2} - \frac{1}{T_1} = \frac{1}{273.15+37.0} - \frac{1}{273.15+25.0} = -1.30·10^{-4}" K-1.

The value of "\text{ln}(\frac{k_1}{k_2})" is:

"\text{ln}(\frac{k_1}{k_2}) = \text{ln}(\frac{0.100}{3.00}) = -3.40" .

Therefore, the activation energy is:

"E_a = 8.314 \text{ (J/(mol K)})·\frac{-3.40}{-1.30·10^{-4} (\text{K}^{-1})}"

"E_a = 2.18\u00b710^5" J/mol, or 218 kJ/mol.

Answer: The value of (1/T2−1/T1) is -1.30·10-4 K-1.